# A Banach *-algebra associated with the action of a group on a topological space

### v1

@@ 3. The ideal intersection property and consequences
The following is the last lemma needed in order to prove the main theorem of this section.
Lemma 3.9.
Suppose that $I$ is a closed ideal such that $I\cap C(X)=\{0\}$. Then, for every $f=\sum_{s\in G}f_{s}\delta^{s}\in I$, we have that $f_{s}$ vanishes on $\text{Aper}(\Sigma)$ for every $s\in G$.
Proof.
Let $f=\sum_{s\in G}f_{s}\delta^{s}\in I$. We want to prove that $f_{s}(x)=0$ for every $x\in\text{Aper}(\Sigma)$ and for every $s\in G$. It is enough to prove that $f_{e}(x)=0$ for every $x\in\text{Aper}(\Sigma)$ as, given that $I$ is an ideal, $f\delta^{s^{-1}}$ belongs to $I$ for every $s\in G$. We fix $x\in\text{Aper}(\Sigma)$.
We consider the quotient $\ell^{1}(\Sigma)/I$. It is a Banach algebra with the following norm:
$\lVert h+I\rVert=\inf_{j\in I}\lVert h+j\rVert\text{ for every }h\in\ell^{1}(\Sigma).$
Let $q:\ell^{1}(\Sigma)\to\ell^{1}(\Sigma)/I$ be the canonical projection, which is clearly a contraction. We have that $I\cap C(X)=\{0\}$, so we can define a norm in $C(X)$ using the norm in the quotient: we define $\lVert\>\cdot\>\rVert_{\infty}^{\prime}=\lVert\>\cdot\>\rVert\circ q$. This norm turns $C(X)$ into a normed algebra, so by the previous theorem we have that $\lVert\>\cdot\>\rVert_{\infty}\leq\lVert\>\cdot\>\rVert^{\prime}$. But, as $q$ is contractive, we also have that $\lVert\>\cdot\>\rVert^{\prime}\leq\lVert\>\cdot\>\rVert_{\infty}$, so the two norms must coincide. We then have that the restriction of $q$ to $C(X)$ is an isometry.
$\lVert f-(f_{e}+\sum_{s\in F}f_{s}\delta^{s})\rVert<\varepsilon.$
We call $b$ to $\sum_{s\in F}f_{s}\delta^{s}$ and $c$ to $f-(f_{e}+b)$, so we can write $f=f_{e}+b+c$, with $\lVert c\rVert<\varepsilon$. As $F$ is finite, we have $F=\{k_{1}\ldots,k_{N}\}$, and as $x$ belongs to $\text{Aper}(\Sigma)$ we have that $\sigma_{k_{i}}(x)\neq x$ for every $i=1\ldots,N$. Using Lemma 3.7 we obtain unimodular functions $\theta_{1},\ldots,\theta_{M}\in C(X)$ and an open neighborhood $U$ of $x$ such that if we write
$f^{\prime}=\frac{1}{M}\sum_{l=1}^{M}\theta_{l}f\overline{\theta_{l}}=\sum_{s\in G}f^{\prime}_{s}\delta^{s}$
[…]
Now, we have that
$\begin{split}\widetilde{f}&=\varphi f^{\prime}=\frac{1}{M}\sum_{l=1}^{M}\varphi\theta_{l}f\overline{\theta_{l}}\\ &=\frac{1}{M}\sum_{l=1}^{M}\varphi\theta_{l}f_{e}\overline{\theta_{l}}+\frac{1}{M}\sum_{l=1}^{M}\varphi\theta_{l}b\overline{\theta_{l}}+\frac{1}{M}\sum_{l=1}^{M}\varphi\theta_{l}c\overline{\theta_{l}}\\ &=\varphi f_{e}+\frac{1}{M}\sum_{l=1}^{M}\varphi\theta_{l}b\overline{\theta_{l}}+\frac{1}{M}\sum_{l=1}^{M}\varphi\theta_{l}c\overline{\theta_{l}},\end{split}$
as we are using that the funtions $\theta_{l}$ are unimodular in order to get the first summand in the last equality. We have that $b=\sum_{s\in F}f_{s}\delta^{s}$, so
$\frac{1}{M}\sum_{l=1}^{M}\varphi\theta_{l}b\overline{\theta_{l}}=\sum_{s\in F}\varphi f_{s}^{\prime}\delta^{s}=\sum_{i=1}^{N}\varphi f_{k_{i}}^{\prime}\delta^{k_{i}}=0,$
given that $\varphi$ is supported in $U$ and the functions $f_{k_{i}}^{\prime}$ vanish in $U$ for every $i=1,\ldots,N$. So, we have that
$\widetilde{f}=\varphi f_{e}+\frac{1}{M}\sum_{l=1}^{M}\varphi\theta_{l}c\overline{\theta_{l}}.$
We set $\widetilde{c}=\frac{1}{M}\sum_{l=1}^{M}\varphi\theta_{l}c\overline{\theta_{l}}$. Then
Now, as $\widetilde{f}$ is in $I$, we have $q(\widetilde{f})=0$. This implies that $q(\varphi f_{e})=-q(\widetilde{c})$. Then
@@
where we use that $q$ is a contraction and its restriction to $C(X)$ is an isometry. So, we get that $|f_{e}(x)|<\varepsilon$. As $\varepsilon$ was arbitrary, we obtain $f_{e}(x)=0$. ∎
Finally, we arrive at the main theorem of this section. We point that equivalence between (1) and (3) in the theorem can be seen as an interpretation for $\ell^{1}(\Sigma)$ of the more general context [BK24, Theorem 5.9], but the proof we include here is following [dJST12, Theorem 4.1], of which this result is a generalization to arbitrary countable (infinite) discrete groups. In order to prove this result we will make use of the family of representations $\{\pi_{x}:\ell^{1}(\Sigma)\to B(H_{x}),x\in X\}$ that we introduced in the preliminaries (see Section 2.2).
@@
    Every closed nonzero ideal of $\ell^{1}(\Sigma)$ has nonzero intersection with $C(X)$.
@@
    Every closed and self-adjoint nonzero ideal of $\ell^{1}(\Sigma)$ has nonzero intersection with $C(X)$.
@@
    $\Sigma$ is topologically free.
@@
It is evident that (1) implies (2). We will prove first that (2) implies (3). Suppose that $\Sigma$ is not topologically free. Then, by Remark 3.3, there exists some $s_{0}\in G\setminus\{e\}$ such that $\text{Fix}(s_{0})$ has nonempty interior. We can then take a nonzero $f\in C(X)$ such that $\text{supp}\left(f\right)$ is contained in $\text{Fix}(s_{0})$. Define $I$ as the closed and self-adjoint ideal of $\ell^{1}(\Sigma)$ generated by $f-f\delta^{s_{0}}$.
Let $x\in X$, and we consider $\pi_{x}:\ell^{1}(\Sigma)\to B(H_{x})$ the representation associated to $x$. We remember that $H_{x}$ is the Hilbert space with orthonormal basis $\{v_{\beta}:\beta\in J_{x}\}$, where $\{s_{\beta}:\beta\in J_{x}\}$ are representatives of $G/G_{x}=\{s_{\beta}G_{x}:\beta\in J_{x}\}$. We will prove that $\pi_{x}$ vanishes at $f-f\delta^{s_{0}},$ which implies that $\pi_{x}$ vanishes on $I$. Let $\beta\in J_{x}$, we have that there are unique $\gamma\in J_{x}$ and $s^{\prime}\in G_{x}$ such that $s_{0}s_{\beta}=s_{\gamma}s^{\prime}\in s_{\gamma}G_{x}$. Then:
$\begin{split}\pi_{x}(f-f\delta^{s_{0}})v_{\beta}&=\pi_{x}(f)v_{\beta}-\pi_{x}(f)\pi_{x}(\delta^{s_{0}})v_{\beta}\\ &=f(\sigma_{s_{\beta}}(x))v_{\beta}-\pi_{x}(f)v_{\gamma}\\ &=f(\sigma_{s_{\beta}}(x))v_{\beta}-f(\sigma_{s_{\gamma}}(x))v_{\gamma}.\end{split}$
Suppose first that $\sigma_{s_{\beta}}(x)\in\text{Fix}(s_{0})$. Then, we have that
@@
and as $\{s_{\beta}:\beta\in J_{x}\}$ are representatives of $G/G_{x}$, it must be that $\beta$ equals $\gamma$. Then it’s clear that $\pi_{x}(f-f\delta^{s_{0}})v_{\beta}=0$. Suppose now that $\sigma_{s_{\beta}}(x)\not\in\text{Fix}(s_{0})$. As $f$ is supported on $\text{Fix}(s_{0})$, we have that $f(\sigma_{s_{\beta}}(x))=0$. We can write $s_{\gamma}=s_{0}s_{\beta}(s^{\prime})^{-1}$. Then
@@
As $\text{Fix}(s_{0})$ is invariant under $\sigma_{s_{0}}$, we have that $\sigma_{s_{0}}(\sigma_{\beta}(x))\not\in\text{Fix}(s_{0})$ and, again, as $f$ is supported on $\text{Fix}(s_{0})$, it must be that $f(\sigma_{s_{0}}(\sigma_{s_{\beta}}(x)))=0.$ It follows that $\pi_{x}(f-f\delta^{s_{0}})v_{\beta}=0$ in both cases. As this stands true for every $\beta\in J_{x}$, we have that $\pi_{x}$ vanishes at $f-f\delta^{s_{0}}$, and as a consequence it also vanishes on $I$.
@@
Let’s prove that (3) implies (1). Let $I$ be a closed ideal such that $I\cap C(X)=\{0\}$. Let $f=\sum_{s\in G}f_{s}\delta^{s}\in I$. Because of Lemma 3.9 we have that $f_{s}$ vanishes on $\text{Aper}(\Sigma)$ for every $s\in G$. As $\Sigma$ is topologically free we have that $\text{Aper}(\Sigma)$ is dense in $X$, so we actually have that $f_{s}=0$ for every $s\in G$, so $f=0$. It follows that $I=\{0\}$. ∎
Theorem 3.10 allows us to relatively easily characterize dynamical properties of $\Sigma$ as analytic-algebraic properties of $\ell^{1}(\Sigma)$, and vice versa. We will do so in the rest of the section. The proofs of many of these results are heavily inspired by the ones in [dJST12] as, given that we have Theorem 3.10 at our disposal, the arguments are particularly similar now.
@@
We say that $\Sigma$ is minimal if there are no proper closed subsets of $X$ that are invariant under the action of $G$. This is equivalent to the fact that the orbit of every element of $X$ is dense in $X$.
Our next objective is to characterize $\Sigma$ being minimal as an analytic-algebraic property of $\ell^{1}(\Sigma)$. In order to do so, we will have to understand how $G$-invariant and closed subsets of $X$ give rise to closed (and self-adjoint) ideals of $\ell^{1}(\Sigma)$.
Let $D\subset X$ be a closed subset invariant under the action of $G$. We can then define the following:
$I_{D}=\left\{f=\sum_{s\in G}f_{s}\delta^{s}\in\ell^{1}(\Sigma):f_{s}|_{D}=0\text{ for every }s\in G\right\}.$
It’s straightforward to check that $I_{D}$ is a closed (and self-adjoint) ideal of $\ell^{1}(\Sigma)$. The following propositions tell that we can go the other way around, that is, to obtain a closed $G$-invariant subset of $X$ from a closed ideal of $\ell^{1}(\Sigma)$.
@@
Let $I$ be a closed ideal of $\ell^{1}(\Sigma)$. Then there exists a closed and $G$-invariant subset $D$ of $X$ such that
@@
As $I$ is a closed ideal of $\ell^{1}(\Sigma)$ it is obvious that $I\cap C(X)$ is a closed ideal of $C(X)$. It is well known that every closed ideal of $\ell^{1}(\Sigma)$ consists of the continuous functions that vanish on some closed subset of $X$. So, there is some closed subset $D$ of $X$ such that
@@
We just have to check that $D$ is invariant under the action of $G$. Suppose that there is some $s\in G$ such that $\sigma_{s}(D)\not\subset D$. This means that there is some $x\in D$ such that $\sigma_{s}(x)\not\in D$. By Urysohn’s lemma we can take $g\in C(X)$ such that $g$ vanishes on $D$ and $g(\sigma_{s}(x))\neq 0$. As $I$ is an ideal, $(\delta^{s})^{-1}g\delta^{s}=g\circ\sigma_{s}\in I$, so in particular it belongs to $I\cap C(X)$. It follows that $g((x))=0$, a contradiction. ∎
A few remarks about the proposition above. First of all, it is not clear if given two different closed ideals $I$ and $J$ of $\ell^{1}(\Sigma)$ we can get the same closed $G$-invariant subset. This has to do with the ability of $C(X)$ to distinguish different ideals through the intersection with it. We will come back to this at the end of this section. In general, it could happen that given a closed ideal of $\ell^{1}(\Sigma)$ we get the whole space $X$ through the proposition above. But when $\Sigma$ is topologically free, Theorem 3.10 tells us that $I\cap C(X)$ is nonzero, so the closed $G$-invariant subset we obtain through the proposition must be a proper subset of $X$.
@@
Suppose that $\Sigma$ is topologically free. Then, the following are equivalent:
@@
    $\ell^{1}(\Sigma)$ has no proper closed ideals.
@@
    $\ell^{1}(\Sigma)$ has no proper closed and self-adjoints ideals.
@@
That (1) implies (2) is evident. Let’s prove that (2) implies (3). Suppose that $\Sigma$ is not minimal, then there is some $x\in X$ such that its orbit is not dense in $X.$ Let $D$ be the closure of the orbit of $x$. Then $D$ is a closed and $G$-invariant proper subset of $X$, so
$I_{D}=\left\{f=\sum_{s\in G}f_{s}\delta^{s}\in\ell^{1}(\Sigma):f_{s}|_{D}=0\text{ for every }s\in G\right\}$
is a proper closed and self-adjoint ideal of $\ell^{1}(\Sigma).$ This contradicts (2).
We will now prove that (3) implies (1). Let $I$ be a closed nonzero ideal of $\ell^{1}(\Sigma)$. As $\Sigma$ is topologically free, Theorem 3.10 tells us that $I\cap C(X)$ is nonzero. So, using Proposition 3.12 we get a closed $G$-invariant subset $D$ that is not the whole space. But as $\Sigma$ is minimal, the only $G$-invariant and closed subsets of $X$ are itself and the empty set. So it must be that $D$ is the empty set, so $I\cap C(X)=C(X)$. This implies that $I=\ell^{1}(\Sigma)$. ∎
@@
We say that $\Sigma$ is topologically transitive if for every pair of nonempty open subsets $U,V$ of $X$ there is some $s\in G$ such that $\sigma_{s}(U)\cap V$ is nonempty.
We will characterize when $\Sigma$ is topologically transitive in terms of the properties of the algebra. We need the following lemma.
@@
    $\Sigma$ is not topologically transitive.
@@
That (1) implies (2) is trivial. Let’s prove that (2) implies (1). As $\Sigma$ is not topologically transitive there are two nonempty and open subsets $U,V$ of $X$ such that $\sigma_{s}(U)\cap V$ is empty for every $s\in G$. We take $O_{1}=\bigcup_{s\in G}\sigma_{s}(U).$ It’s clear that $O_{1}$ is an open $G$-invariant subset of $X$. As $V\subset\overline{O_{1}}^{c}$, we can take $O_{2}=\overline{O_{1}}^{c}$. These sets verify what we need. ∎
The following theorem generalizes [dJST12, Theorem 4.10], which was proved for $\mathbb{Z}$-actions. It characterizes some notion of primality of the algebra as the action being topologically transitive. In the proof, given a closed subset $D$ of $X$, we will denote $\ker(D)=\left\{g\in C(X):g|_{D}=0\right\}.$ We recall that $E$ denotes the the canonical projection from $\ell^{1}(\Sigma)$ to $C(X)$ given in Definition 2.1.
@@
Suppose that $\Sigma$ is topologically free. Then, the following are equivalent:
@@
    For every pair of nonzero closed ideals $I,J$ of $\ell^{1}(\Sigma)$, we have that $I\cap J$ is nonzero.
@@
    For every pair of nonzero closed and self-adjoints ideals $I,J$ of $\ell^{1}(\Sigma)$, we have that $I\cap J$ is nonzero.
@@
    $\Sigma$ is topologically transitive, i.e, for every pair of open non-empty subsets $U,V$ of $X$, there is $s\in G$ such that $V$ intersects the image of $U$ through the action of $s$.
@@
It is evident that (1) implies (2). Let’s see that (2) implies (3). Suppose that $\Sigma$ is not topologically transitive, then by Lemma 3.15 we have two nonempty open sets $O_{1}$ and $O_{2}$, disjoint and $G$-invariant, such that $X=\overline{O_{1}}\cup\overline{O_{2}}$. We define
$I_{i}=\left\{\sum_{s\in G}f_{s}\delta^{s}:f_{s}\in\ker(\overline{O_{i}})\text{for every }s\in G\right\}$
for $i=1,2$. Both are closed and self-adjoint ideals of $\ell^{1}(\Sigma)$. Then
@@
Let’s now prove that (3) implies (1). Let $I_{1}$ and $I_{2}$ be two proper and closed ideals of $\ell^{1}(\Sigma)$ such that $I_{1}\cap I_{2}=\{0\}$. As $\Sigma$ is topologically free, Theorem 3.10 tells us that $I_{1}\cap C(X)$ and $I_{2}\cap C(X)$ are both proper ideals of $C(X)$, so by Proposistion 3.12 we get two proper closed and $G$-invariant subsets $D_{1}$ and $D_{2}$ of $X$ such that $I_{1}\cap C(X)=\ker(D_{1})$ and $I_{2}\cap C(X)=\ker(D_{2})$. As $I_{1}\cap I_{2}=\{0\}$ we also have that $I_{1}\cap I_{2}\cap C(X)=\{0\}$. Then
@@
so it must be $D_{1}\cup D_{2}=X$. As both are proper and closed, they must both have nonempty interior, so if we denote $U$ the interior of $D_{1}$ and $V$ the complement of $D_{1}$ we have that $\sigma_{s}(U)\cap V=\emptyset$ for every $s\in G$. This contradicts the topological transitivity of $\Sigma$. ∎
In Proposition 3.12 we mentioned a way of relating closed ideals of $\ell^{1}(\Sigma)$ to closed ideals of $C(X)$: if $I$ is a closed ideal of $\ell^{1}(\Sigma)$, then $I\cap C(X)$ is a closed ideal of $C(X)$. A reasonable question to ask is whether $C(X)$ detects all closed ideals of $\ell^{1}(\Sigma)$ through this procedure, i.e., if $I$ is a closed nonzero ideal of $\ell^{1}(\Sigma)$, then $I\cap C(X)$ is a nonzero ideal of $C(X)$. This property is a case of the general property called the ideal intersection property. This property has been studied in different contexts, see, for example, [AT25, AR23], where the property was studied between different (reduced) $C^{*}$-algebras and subalgebras of them. In our context, Theorem 3.10 tells us exactly when this holds, that is, when the action is topologically free. Another similar property is the ideal separation property, which was also studied in [AR23]. This property holds if given two closed ideals $I$ and $J$ of $\ell^{1}(\Sigma)$, $I\cap C(X)=J\cap C(X)$ implies $I=J$. So, this property holds if we can distinguish closed ideals of $\ell^{1}(\Sigma)$ through their intersection with $C(X)$, and Theorem 3.10 tells us that for this to hold we need at least the action to be topologically free. It turns out that this ideal separation property is related to the definition that follows. This definition appeared before, for example, in [Sie10, Definition 1.13] and [BK24, Definition 6.11].

### v2

@@ 3. The ideal intersection property and consequences
As $\varepsilon$ was arbitrary, we conclude that $f_{e}(x)=0$. ∎
Finally, we arrive at the main theorem of this section. We point that equivalence between (1) and (3) in the theorem can be seen as an interpretation for $\ell^{1}(G\curvearrowright X)$ of the more general context [BK24, Theorem 5.9], but the proof we include here is following [dJST12, Theorem 4.1], of which this result is a generalization to arbitrary countable (infinite) discrete groups. In order to prove this result we make use of the family of representations $\{\pi_{x}:\ell^{1}(G\curvearrowright X)\to B(H_{x}),x\in X\}$ that we introduced in the preliminaries (see Section 2.2).
@@
    For every closed nonzero ideal $I$ of $\ell^{1}(G\curvearrowright X)$ the intersection $I\cap C(X)$ is nonzero.
@@
    For every closed and self-adjoint nonzero ideal of $\ell^{1}(G\curvearrowright X)$ the intersection $I\cap C(X)$ is nonzero.
@@
    The action $G\curvearrowright X$ is topologically free.
@@
It is clear that (1) implies (2). We prove first that (2) implies (3). Suppose that the action $G\curvearrowright X$ is not topologically free. Then, by Remark 3.3, there exists some $s_{0}\in G\setminus\{e\}$ such that $\text{Fix}(s_{0})$ has nonempty interior. We can then take a nonzero $f\in C(X)$ such that $\text{supp}\left(f\right)$ is contained in $\text{Fix}(s_{0})$. We define $I$ as the closed and self-adjoint ideal of $\ell^{1}(G\curvearrowright X)$ generated by $f-f\delta_{s_{0}}$.
Let $x\in X$, and we consider $\pi_{x}:\ell^{1}(G\curvearrowright X)\to B(H_{x})$ the representation associated to $x$. We remember that $H_{x}$ is the Hilbert space with orthonormal basis $\{v_{\beta}:\beta\in J_{x}\}$, where $\{s_{\beta}:\beta\in J_{x}\}$ is a set of representatives of the quotient $G/G_{x}=\{s_{\beta}G_{x}:\beta\in J_{x}\}$. We will prove that $\pi_{x}$ vanishes at $f-f\delta_{s_{0}},$ which implies that $\pi_{x}$ vanishes on $I$. Let $\beta\in J_{x}$, we have that there are unique $\gamma\in J_{x}$ and $s^{\prime}\in G_{x}$ such that $s_{0}s_{\beta}=s_{\gamma}s^{\prime}\in s_{\gamma}G_{x}$. Then:
$\begin{split}\pi_{x}(f-f\delta_{s_{0}})v_{\beta}&=\pi_{x}(f)v_{\beta}-\pi_{x}(f)\pi_{x}(\delta_{s_{0}})v_{\beta}=f(\sigma_{s_{\beta}}(x))v_{\beta}-\pi_{x}(f)v_{\gamma}\\ &=f(\sigma_{s_{\beta}}(x))v_{\beta}-f(\sigma_{s_{\gamma}}(x))v_{\gamma}.\end{split}$
Suppose first that $\sigma_{s_{\beta}}(x)$ is in $\text{Fix}(s_{0})$. Then, we have that
@@
and as $\{s_{\beta}:\beta\in J_{x}\}$ are representatives of $G/G_{x}$, it must be that $\beta$ equals $\gamma$. Then, it’s clear that $\pi_{x}(f-f\delta_{s_{0}})v_{\beta}=0$. Suppose now that $\sigma_{s_{\beta}}(x)$ is not in $\text{Fix}(s_{0})$. As $f$ is supported on $\text{Fix}(s_{0})$, we have that $f(\sigma_{s_{\beta}}(x))=0$. We can write $s_{\gamma}=s_{0}s_{\beta}(s^{\prime})^{-1}$. Then
@@
As $\text{Fix}(s_{0})$ is invariant under $\sigma_{s_{0}}$, we have that $\sigma_{s_{0}}(\sigma_{\beta}(x))$ is not in $\text{Fix}(s_{0})$ and, again, as $f$ is supported on $\text{Fix}(s_{0})$, it must be that $f(\sigma_{s_{0}}(\sigma_{s_{\beta}}(x)))=0.$ It follows that $\pi_{x}(f-f\delta_{s_{0}})v_{\beta}=0$ in both cases. As this stands true for every $\beta\in J_{x}$, we have that $\pi_{x}$ vanishes at $f-f\delta_{s_{0}}$, and as a consequence it also vanishes on $I$.
@@
Now, let’s prove that (3) implies (1). Let $I$ be a closed ideal of $\ell^{1}(G\curvearrowright X)$ such that $I\cap C(X)=\{0\}$. Let $f=\sum_{s\in G}f_{s}\delta_{s}\in I$. Because of Lemma 3.6 we have that $f_{s}$ vanishes on $\text{Aper}(G\curvearrowright X)$, for every $s\in G$. As $G\curvearrowright X$ is topologically free we have that $\text{Aper}(G\curvearrowright X)$ is dense in $X$, it follows that $f_{s}=0$ for every $s\in G$, so $f=0$. This implies that $I=\{0\}$. ∎
Theorem 3.7 allows us to relatively easily characterize dynamical properties of $G\curvearrowright X$ as analytic-algebraic properties of $\ell^{1}(G\curvearrowright X)$, and vice versa. We do so in the rest of this section. The proofs of many of these results are heavily inspired by the ones in [dJST12] as, given that we have Theorem 3.7 at our disposal, the arguments are particularly similar now.
@@
We say that the action $G\curvearrowright X$ is minimal if there are no proper closed subsets of $X$ that are invariant under the action of $G$. This is equivalent to the orbit of every element being dense in $X$.
Our next objective is to characterize the action being minimal as an analytic-algebraic property of $\ell^{1}(G\curvearrowright X)$. In order to do so, we have to understand how $G$-invariant and closed subsets of $X$ give rise to closed (and self-adjoint) ideals of $\ell^{1}(G\curvearrowright X)$.
Let $D$ be a closed subset of $X$ invariant under the action of $G$. We can define the following:
$I_{D}=\left\{f=\sum_{s\in G}f_{s}\delta_{s}\in\ell^{1}(G\curvearrowright X):f_{s}|_{D}=0\text{ for every }s\in G\right\}.$
It’s straightforward to check that $I_{D}$ is a closed and self-adjoint ideal of $\ell^{1}(G\curvearrowright X)$. The following propositions tell that we can go the other way around, that is, to obtain a closed $G$-invariant subset of $X$ from a closed ideal of $\ell^{1}(G\curvearrowright X)$.
@@
Let $I$ be a closed ideal of $\ell^{1}(G\curvearrowright X)$. Then, there exists a closed and $G$-invariant subset $D$ of $X$ such that
@@
As $I$ is a closed ideal of $\ell^{1}(G\curvearrowright X)$ it is clear that $I\cap C(X)$ is a closed ideal of $C(X)$. It is well known that every closed ideal of $\ell^{1}(G\curvearrowright X)$ consists of the continuous functions that vanish on some closed subset of $X$. This means that there is some closed subset $D$ of $X$ such that
@@
We just have to check that $D$ is invariant under the action of $G$. Suppose that there is some $s\in G$ such that $\sigma_{s}(D)$ is not included in $D$. This means that there is some $x\in D$ such that $\sigma_{s}(x)\not\in D$. By Urysohn’s lemma we can take $g\in C(X)$ such that $g$ vanishes on $D$ and $g(\sigma_{s}(x))\neq 0$. As $I$ is an ideal we have that $(\delta_{s})^{-1}g\delta_{s}=g\circ\sigma_{s}\in I$, so in particular it belongs to $I\cap C(X)$. It follows that $g(x)=0$, which is a contradiction. ∎
@@
Suppose that the action $G\curvearrowright X$ is topologically free. Then, the following are equivalent:
@@
    There are no proper closed ideals in $\ell^{1}(G\curvearrowright X)$.
@@
    There are no proper closed and self-adjoint ideals in $\ell^{1}(G\curvearrowright X)$.
@@
That (1) implies (2) is evident. Let’s prove that (2) implies (3). Suppose that the action $G\curvearrowright X$ is not minimal, then there is some $x\in X$ such that its orbit is not dense in $X.$ Let $D$ be the closure of the orbit of $x$. Then $D$ is a closed and $G$-invariant proper subset of $X$, so
$I_{D}=\left\{f=\sum_{s\in G}f_{s}\delta_{s}\in\ell^{1}(G\curvearrowright X):f_{s}|_{D}=0\text{ for every }s\in G\right\}$
is a proper closed and self-adjoint ideal of $\ell^{1}(G\curvearrowright X).$ This contradicts (2).
We now prove that (3) implies (1). Let $I$ be a closed nonzero ideal of $\ell^{1}(G\curvearrowright X)$. As the action $G\curvearrowright X$ is topologically free, Theorem 3.7 tells us that $I\cap C(X)$ is nonzero. So, using Proposition 3.9 we get a closed $G$-invariant subset $D$ that is not the whole space. But as $G\curvearrowright X$ is minimal, the only $G$-invariant and closed subsets of $X$ are itself and the empty set. Then, it must be that $D$ is the empty set, so $I\cap C(X)=C(X)$. This implies that $I=\ell^{1}(G\curvearrowright X)$. ∎
@@
We say that the action $G\curvearrowright X$ is topologically transitive if for every pair of nonempty open subsets $U$ and $V$ of $X$ there is some $s\in G$ such that $\sigma_{s}(U)\cap V$ is nonempty.
We characterize when the action is topologically transitive in terms of the properties of the algebra. We need the following lemma.
@@
    The action $G\curvearrowright X$ is not topologically transitive.
@@
That (1) implies (2) is trivial. Let’s prove that (2) implies (1). As $G\curvearrowright X$ is not topologically transitive, there are two nonempty and open subsets $U$ and $V$ of $X$ such that $\sigma_{s}(U)\cap V$ is empty for every $s\in G$. Set $O_{1}=\bigcup_{s\in G}\sigma_{s}(U).$ It’s clear that $O_{1}$ is an open $G$-invariant subset of $X$. As $V$ is contained in $\overline{O_{1}}^{c}$, we can take $O_{2}=\overline{O_{1}}^{c}$. These sets verify what we need. ∎
The following theorem generalizes [dJST12, Theorem 4.10], which was proved for $\mathbb{Z}$-actions. It characterizes some notion of primality of the algebra as the action being topologically transitive. In the proof, given a closed subset $D$ of $X$, we denote by $\ker(D)$ the set $\left\{g\in C(X):g|_{D}=0\right\}.$ We recall that $E$ denotes the the canonical projection from $\ell^{1}(G\curvearrowright X)$ to $C(X)$ given in Definition 2.1.
@@
Suppose that the action $G\curvearrowright X$ is topologically free. Then, the following are equivalent:
@@
    If $I$ and $J$ are closed nonzero ideals of $\ell^{1}(G\curvearrowright X)$ then $I\cap J$ is nonzero.
@@
    If $I$ and $J$ are closed and self-adjoint nonzero ideals of $\ell^{1}(G\curvearrowright X)$ then $I\cap J$ is nonzero.
@@
    The action $G\curvearrowright X$ is topologically transitive.
@@
It is evident that (1) implies (2). Let’s see that (2) implies (3). Suppose that the action $G\curvearrowright X$ is not topologically transitive, then by Lemma 3.12 we have two nonempty open sets $O_{1}$ and $O_{2}$, disjoint and $G$-invariant, such that $X=\overline{O_{1}}\cup\overline{O_{2}}$. We define
$I_{i}=\left\{\sum_{s\in G}f_{s}\delta_{s}:f_{s}\in\ker(\overline{O_{i}})\text{ for every }s\in G\right\}$
for $i=1,2$. Both are closed and self-adjoint ideals of $\ell^{1}(G\curvearrowright X)$. Then
@@
Let’s now prove that (3) implies (1). Let $I_{1}$ and $I_{2}$ be two proper and closed ideals of $\ell^{1}(G\curvearrowright X)$ such that $I_{1}\cap I_{2}=\{0\}$. As $G\curvearrowright X$ is topologically free, Theorem 3.7 tells us that $I_{1}\cap C(X)$ and $I_{2}\cap C(X)$ are both proper ideals of $C(X)$, so by Proposistion 3.9 we get two proper closed and $G$-invariant subsets $D_{1}$ and $D_{2}$ of $X$ such that $I_{1}\cap C(X)=\ker(D_{1})$ and $I_{2}\cap C(X)=\ker(D_{2})$. As $I_{1}\cap I_{2}=\{0\}$ we also have that $I_{1}\cap I_{2}\cap C(X)=\{0\}$. Then
@@
so it must be $D_{1}\cup D_{2}=X$. As both are proper and closed, they must both have nonempty interior, so if we denote by $U$ the interior of $D_{1}$ and by $V$ the complement of $D_{1}$ we have that $\sigma_{s}(U)\cap V=\emptyset$ for every $s\in G$. This contradicts the topological transitivity of $G\curvearrowright X$. ∎
In Proposition 3.9 we mentioned a way of relating closed ideals of $\ell^{1}(G\curvearrowright X)$ to closed ideals of $C(X)$: if $I$ is a closed ideal of $\ell^{1}(G\curvearrowright X)$, then $I\cap C(X)$ is a closed ideal of $C(X)$. A reasonable question to ask is whether $C(X)$ detects all closed ideals of $\ell^{1}(G\curvearrowright X)$ through this procedure, i.e., if $I$ is a closed nonzero ideal of $\ell^{1}(G\curvearrowright X)$, then $I\cap C(X)$ is a nonzero ideal of $C(X)$. This property is a case of the general property called the ideal intersection property. This property has been studied in different contexts, see, for example, [AT25, AR23], where the property was studied between different (reduced) $C^{*}$-algebras and subalgebras of them, and [Flo25, Theorem 5.6], where it was studied for the algebra of crossed sections associated to a Fell bundle over a group with polynomial growth and its universal $C^{*}$-algebra completion. In our context, Theorem 3.7 tells us exactly when this holds between $\ell^{1}(G\curvearrowright X)$ and $C(X)$, that is, when the action is topologically free.
Another similar property is the ideal separation property, which was also studied in [AR23]. This property holds if, for any pair of closed ideals of $\ell^{1}(G\curvearrowright X)$, they are equal if they have the same intersection with $C(X)$. Explicitly, $I\cap C(X)=J\cap C(X)$ implies $I=J$ for every pair of closed ideals $I$ and $J$ of $\ell^{1}(G\curvearrowright X)$. So, this property holds if we can distinguish closed ideals of $\ell^{1}(G\curvearrowright X)$ through their intersection with $C(X)$, and Theorem 3.7 tells us that for this to hold we need at least the action to be topologically free. It turns out that the ideal separation property is related to the definition that follows, which is a strengthening of topological freeness.